OpenStudy (anonymous):

y=-x^2 + 2x + 8 What is the range

OpenStudy (anonymous):

you need help?

OpenStudy (anonymous):

the range is set of values that y can assume given a value of x. if we think about this by first picturing the graph in our head.. that would be best.

OpenStudy (anonymous):

I wanted to know the math steps to do it

OpenStudy (amoodarya):

y=-x^2 + 2x + 8 y=-(x^2 - 2x) + 8 y=-((x-1)^2-1)) + 8 y=-(x-1)^2+1+8 y=-(x-1)^2+9 now y-9=-(x-1)^2 9-y=(x-1)^2 >=0 so 9-y>=0 so y<=9

OpenStudy (anonymous):

ok sorry. i just find it easier to think of this as the graph of -x^2 shifted up by 8 units and right by 1.

OpenStudy (amoodarya):

"Nameless" it is wrong because by shifting +1 right and shif +8 range became y<8 but another method is using discriminant x^2-2x-8+y=0 delta >=0 4-4(-8+y)>=0 1+8 -y>=0 y<=9

OpenStudy (anonymous):

ahh i totally overlooked that. nice.

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

amoodarya can you please explain were you got the 4-4 in front of (-8+y)